OpenStudy (anonymous):

Will award medal Find the solution to the differential equation (dB/dt)+4B=10 B(1)=90

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

\[\frac{dB}{dt}+4B=10 \\ \frac{dB}{dt}=10-4B \\ \frac{dB}{10-4B}=dt \] integrate both sides

myininaya (myininaya):

if the left hand side looks ugly use a substitution

OpenStudy (anonymous):

i can prolly integrate it from then i just was confused by the seperation of variables

myininaya (myininaya):

Oh well do you understand how I separated it?

OpenStudy (anonymous):

oh yea ive done like 20 of these

myininaya (myininaya):

oh ok well let me know if I can help anymore i will be around a little longer

OpenStudy (anonymous):

so you get -4B=B(e^t)-10

myininaya (myininaya):

what did you on both sides before doing anything else

myininaya (myininaya):

like directly after the integration

OpenStudy (anonymous):

oh you have ln|10-4b|=t+c ----> |10-4b|=e^(t+c)

myininaya (myininaya):

you are missing something in front of the ln part

myininaya (myininaya):

u=10-4b du=-4 dB so -1/4 du=dB

myininaya (myininaya):

so we should have \[\frac{-1}{4} \ln|10-4b|=t+c\]

myininaya (myininaya):

then you can multiply the -4 on both sides then do the whole e^x and ln(x) are inverse thing

OpenStudy (anonymous):

B=-350e^(-4t)-10

myininaya (myininaya):

oops and wherever I had little b I meant big B :p

OpenStudy (anonymous):

that shouldnt change much

OpenStudy (anonymous):

i mean at least for my answer

myininaya (myininaya):

lol no not really you have some really big numbers for your B

myininaya (myininaya):

oh you apply the condition

OpenStudy (anonymous):

yea...

myininaya (myininaya):

hmm...I will show you what I did

myininaya (myininaya):

\[\frac{-1}{4} \ln|10-4B|=t+C \\ \ln|10-4B|=-4t+C \\ 10-4B=e^{-4t+C} \\ -4B=e^{-4t+C}-10 \\ B=\frac{-1}{4} e^{-4t +C}+\frac{10}{4} \\ B=\frac{-K}{4} e^{-4t}+\frac{5}{2} \\ B=K_1e^{-4 t}+\frac{5}{2} \\ \text{ so applying the } B(1)=90 \\ 90=K_1e^{-4}+\frac{5}{2} \\ 90-\frac{5}{2}=K_1e^{-4} \\ \frac{175}{2}=K_1e^{-4} \\ \frac{175}{2}e^4=K_1\]

myininaya (myininaya):

like in the first step I could have written -4C but that is still a constant and that is also why I recalled -k/4 another constant value

myininaya (myininaya):

it doesn't matter if we multiply the unknown constant by another constant it is still the unknown constant that we must find

OpenStudy (anonymous):

the unknown constant is -362.5

myininaya (myininaya):

that depends on what you used as the equation to find the unknown constant as

myininaya (myininaya):

the thing is ending up with the same equation

OpenStudy (anonymous):

90=b(-1/4)(e^4t)+5/2

myininaya (myininaya):

\[B=K_1e^{-4t}+\frac{5}{2} \\ \text{ where } K_1=\frac{175}{2}e^4\]

myininaya (myininaya):

is b your constant?

OpenStudy (anonymous):

yea

myininaya (myininaya):

\[90=\frac{-b}{4}e^{4}+\frac{5}{2} \\ -360=be^{4}+10 \\ -350=be^{4} \\ b=\frac{-350}{e^4}\] ok so you would have \[B=\frac{350}{4e^4}e^{-4t}+\frac{5}{2} \] but I believe 350/4 reduced to 175/2 :)

myininaya (myininaya):

same exact equation just a a bit of different name for the constant we were finding

myininaya (myininaya):

oops wait I think there was a negative missing in your equation one sec

OpenStudy (anonymous):

yea i tried it like you did and my webwork was wrong

myininaya (myininaya):

90=b(-1/4)(e^4t)+5/2 this was B=b(-1/4)e^(-4t)+5/2 right

myininaya (myininaya):

the first way was right

myininaya (myininaya):

but not the one with this newer constant

myininaya (myininaya):

anyways it was suppose to be B=b(-1/4)e^(-4t)+5/2

myininaya (myininaya):

so apply your B(1)=90

OpenStudy (anonymous):

alright ill give it a whirl

myininaya (myininaya):

\[90=b(\frac{-1}{4})e^{-4}+\frac{5}{2} \\ -360=be^{-4}-10 \\ -350=be^{-4} \\ -350e^4=b\]

myininaya (myininaya):

the e^4 was suppose to be on top not on bottom

myininaya (myininaya):

\[B=b (\frac{-1}{4})e^{-4t}+\frac{5}{2} \\ \text{ where } b=-350e^4 \]

OpenStudy (anonymous):

(-362.5e^4)/4e^(-4)+(5/2)

myininaya (myininaya):

well almost

myininaya (myininaya):

neg * neg =pos

myininaya (myininaya):

and you are missing the t in the ^(-4) thing

myininaya (myininaya):

and where does 362.5 come from?

myininaya (myininaya):

continuing from what I had earlier \[\frac{-1}{4} \ln|10-4B|=t+C \\ \ln|10-4B|=-4t+C \\ 10-4B=e^{-4t+C} \\ -4B=e^{-4t+C}-10 \\ B=\frac{-1}{4} e^{-4t +C}+\frac{10}{4} \\ B=\frac{-K}{4} e^{-4t}+\frac{5}{2} \\ B=K_1e^{-4 t}+\frac{5}{2} \\ \text{ so applying the } B(1)=90 \\ 90=K_1e^{-4}+\frac{5}{2} \\ 90-\frac{5}{2}=K_1e^{-4} \\ \frac{175}{2}=K_1e^{-4} \\ \frac{175}{2}e^4=K_1\] \[B=K_1e^{-4t}+\frac{5}{2} \\ \text{ where } K_1=\frac{175}{2}e^{4} \\ \\ B=\frac{175}{2}e^4e^{-4t}+\frac{5}{2} \\ B=\frac{175}{2}e^{4-4t}+\frac{5}{2} \\ B=\frac{5}{2}(35e^{4-4t}+1)\] This is the same result you will get using your equation and your constant ... \[B=b (\frac{-1}{4})e^{-4t}+\frac{5}{2} \\ \text{ where } b=-350e^4 \\ B=-350e^4(\frac{-1}{4})e^{-4t}+\frac{5}{2} \\ B=\frac{350}{4}e^{4-4t}+\frac{5}{2} \\ B=\frac{175}{2}e^{4-4t}+\frac{5}{2} \\ B=\frac{5}{2}(35e^{4-4t}+1)\]

myininaya (myininaya):

I know the algebra looks scary

myininaya (myininaya):

You just have to take your time and work through it

myininaya (myininaya):

Please tell me if you don't understand something up . Or if you think I need to work through each step one at a time with you.

myininaya (myininaya):

@pmkat14 If you are still there, I'm going to leave in a few. Just letting you know if you had any question.